if one person can do a job in 8 hours, and two people working together can do it in 6 hours, how long would it take the second person to do it alone?
Answer by oberobic(2304)

The key to solving work problems like this is to determine how much of the job each per can do per unit of time. Then you solve how long it will take to total 1 job.
.
One person can do the job in 8 hr, so he does 1/8 of the job per hour.
Two people working together can do the job in 6 hr, so their combined speed is 1/6 of the job per hour.
.
1/8 + 1/x = 1/6
1/x = 1/6 - 1/8
1/x = 2/48 = 1/24
So, the second person takes 24 hr to do the job.
.
Checking how much they get done in 6 hours working together.
6 hr * 1/8 job/hr = 3/4 of the job
6 hr * 1/24 job/hr = 1/4 of the job
3/4 job + 1/4 job = 1 job
.
Answer:
The second person can do the job in 24 hr working alone.
.
Done